(no subject)

arenton@uvahee.phys.virginia.edu
Sun, 19 Dec 1999 16:46:32 -0500

Shift Summary for day shift, 19/dec/99

On Shift:Bob T, Mike A

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:power outage recovery continues

========================================================================

Tapes brought to FCC this shift: UPA,B,C 232-237,UPS 78,79 plus
various onsplit.

========================================================================
Accesses during shift: open

========================================================================
Overview of shift:

At 0800 the estimate appears that beam will return Tuesday.

Did full laser scan.

Tried a cosmic run with tapes, it works.

Bob T. went to 13th floor and rebooted vivsgi, so these
messages now get somewhere.